OpenStudy (anonymous):

Solve for x @confluxepic -6x+14=12-8x

OpenStudy (anonymous):

@confluxepic

OpenStudy (confluxepic):

-6x + 14 = 12 - 8x (+8x) (+8x) 2x + 14 = 12 2x = -2 x = -1

OpenStudy (anonymous):

thanks.

OpenStudy (confluxepic):

This could help you get started in Open Study.